solve for x
3x + 1/2 = 1/4

Answer by ankor@dixie-net.com(22740)

solve for x
3x + 1/2 = 1/4
:
multiply equation by 4 to get rid of the denominators:
12x + 2 = 1
:
12x= 1 - 2
:
x = -1/12
:
Check for x= -1/12 in the original equation:
3(-1/12) + 1/2 =
-3/12 + 1/2 =
-3/12 + 6/12 = +3/12 which is 1/4